# Scalable In-Context Q-Learning

### v1

@@ Abstract
Recent advancements in language models have demonstrated remarkable in-context learning abilities, prompting the exploration of in-context reinforcement learning (ICRL) to extend the promise to decision domains. Due to involving more complex dynamics and temporal correlations, existing ICRL approaches may face challenges in learning from suboptimal trajectories and achieving precise in-context inference. In the paper, we propose Scalable In-Context Q-Learning (SICQL), an innovative framework that harnesses dynamic programming and world modeling to steer ICRL toward efficient reward maximization and task generalization, while retaining the scalability and stability of supervised pretraining. We design a prompt-based multi-head transformer architecture that simultaneously predicts optimal policies and in-context value functions using separate heads. We pretrain a generalized world model to capture task-relevant information, enabling the construction of a compact prompt that facilitates fast and precise in-context inference. During training, we perform iterative policy improvement by fitting a state value function to an upper-expectile of the Q-function, and distill the in-context value functions into policy extraction using advantage-weighted regression. Extensive experiments across a range of discrete and continuous environments show consistent performance gains over various types of baselines, especially when learning from suboptimal data. Our code is available at https://github.com/NJU-RL/SICQL.
@@ 1 Introduction
A longstanding goal of reinforcement learning (RL) is to learn from diverse experiences and generalize beyond its training environments, efficiently adapting to unseen situations, dynamics, or objectives (ackley1989generalization, ; kirk2023survey, ). Recently, large language models (LLMs) (achiam2023gpt, ; guo2025deepseek, ) have exhibited tremendous success in training large-scale transformer models on massive datasets to achieve remarkable in-context learning abilities, i.e., adapting to new tasks via prompt conditioning without any model updates (brown2020language, ; li2024visual, ; wang2023context, ). Accordingly, in-context RL (ICRL) seeks to extend this promise to decision domains and has seen rapid progress in recent years (nikulin2025xland, ). Existing studies in offline settings contain two typical branches: algorithm distillation (AD) (laskin2022context, ) and decision-pretrained transformer (DPT) (lee2023supervised, ), due to their simplicity and generality. They commonly employ cross-episode transitions as few-shot prompts and train transformer-based policies under supervised pretraining (lin2024transformers, ), followed by various improvements from model-based planning (son2025distilling, ), hierarchical decomposition (huang2024context, ), importance weighting (dong2024incontext, ), etc (sinii2024context, ; tarasov2025yes, ; dai2024context, ).
@@
Though, significant challenges may emerge when extending the promise of in-context learning from (self-) supervised learning to RL, since RL involves more complex dynamics and temporal correlations (silver2021reward, ). First, previous studies usually adopt a supervised pretraining paradigm, failing to utilize dynamic programming  [^1] In this paper, we use dynamic programming to indicate the fundamental characteristic of any RL algorithm relying on the Bellman-backup operation. It updates state (or state-action) values based on value estimates of successor states (or state-action pairs), i.e., bootstrapping (sutton1998reinforcement, ). We use dynamic programming and Q-learning interchangeably to refer to this fundamental property of RL. offered by full RL algorithms to go beyond merely leveraging collected data (yamagata2023q, ). AD distills the entire learning history into a causal transformer, which can require long-horizon context and inherit suboptimal behaviors due to the gradual update rule (son2025distilling, ). DPT trains a transformer model to predict the optimal action given a query state and a prompt of interaction transitions, which relies on an oracle for optimal action labeling that can often be infeasible in practice (tarasov2025yes, ). Second, ICRL generally takes raw transitions as prompts that have many more tokens than a sentence and can be highly redundant.  [^2] In language communities, the prompts are concise and precise text instructions with rich semantic information, as texts or words naturally involve high-level concepts that describe entire semantic entities such as objects (nouns) and actions (verbs) (kakogeorgiou2022hide, ). Moreover, within transitions from offline datasets, the task information can be entangled with behavior policies, thus producing biased task inference at test time (yuan2022robust, ). Hence, this type of prompt may be insufficient to precisely capture relevant information about decision tasks. The aforementioned limitations raise a key question: Can we design a scalable ICRL framework using lightweight prompts that precisely capture task-relevant information, while unleashing the core potential of fundamental reward maximization to learn from suboptimal data?
@@
To tackle these challenges, we draw upon two basic properties inherent in full RL. First, classical RL algorithms learn a value function to backpropagate expected returns using dynamic programming updates, showing high stitching capacity, i.e., the ability to combine parts of suboptimal trajectories for finding globally optimal behaviors (sutton1998reinforcement, ). Naturally, the stitching property offers a compelling avenue for unleashing the potential of ICRL architectures, attaining substantial improvement over suboptimal data. Second, RL agents learn through active interactions with the outer world (environment) based on the Markov decision process (MDP) formulation. The world model (environment dynamics) (hafner2025mastering, ) completely characterizes the underlying task, and is intrinsically invariant to behavior policies or collected datasets (wang2024meta, ). Hence, leveraging the world model holds promise for designing a lightweight prompt structure capable of precisely encoding task-relevant information.
@@
Drawing inspiration from full RL, we propose Scalable In-Context Q-Learning (SICQL), an innovative framework that harnesses dynamic programming and world modeling to steer ICRL toward efficient reward maximization and task generalization. First, we design a prompt-based multi-head transformer architecture to maintain scalability and parameter efficiency. The model simultaneously predicts optimal policies and in-context value functions using separate heads, given a task prompt and corresponding query inputs. Second, we pretrain a generalized world model to capture task-relevant information from the multi-task offline dataset, and use it to transform a small number of raw transitions into a lightweight prompt for fast and precise in-context inference. Finally, we perform iterative policy improvement by fitting a state value function to an upper-expectile of the Q-function, and distill the in-context value functions into policy extraction using advantage-weighted regression. This formulation allows for learning a policy to maximize the Q-values subject to an offline dataset constraint, while retaining the scalability and stability of the supervised pretraining paradigm.
In summary, our main contributions are threefold:
@@ 2 Related Work
The concept of agents adapting their behaviors within the context without model updates builds on earlier work in meta-RL, notably RL2 (duan2016rl, ) that includes past interactions as inputs to an RNN-based policy. Recently, there has been a shift towards employing transformers to implement ICRL (nikulin2025xland, ), driven by their proven ability to capture long-term dependencies and exhibit emergent in-context learning behaviors brown2020language . In online RL settings, AMAGO (grigsby2023amago, ; grigsby2024amago2, ) trains long-sequence transformers over entire rollouts with actor-critic learning to tackle in-context goal-conditioned problems, and (lu2023structured, ) leverages the S4 (structured state space sequence) model’s ability to handle long-range sequences for ICRL tasks. In offline settings, two classical branches are AD (laskin2022context, ) and DPT (lee2023supervised, ). AD trains a causal transformer to autoregressively predict actions using preceding learning histories as context, while DPT predicts the optimal action based on a query state and a prompt of interaction transitions. Subsequent studies further increase their in-context learning capacities from various perspectives (son2025distilling, ; huang2024context, ; sinii2024context, ; dai2024context, ; zisman2024emergence, ). For example, IDT (huang2024context, ) designs a hierarchical structure of human decision-making, and DICP (son2025distilling, ) incorporates model-based planning with a learned dynamics model. These approaches generally adopt a supervised paradigm and use raw transitions as prompts, while we harness dynamic programming for reward maximization and leverage world modeling to construct more efficient prompts.
@@
DIT (dong2024incontext, ) and IC-IQL (tarasov2025yes, ) are the most relevant to our work. DIT (dong2024incontext, ) also uses a weighted maximum likelihood estimation loss to train the transformer policy, where the weights are directly calculated from observed rewards in collected datasets. In contrast, we learn separate in-context value functions for computing advantage weights, enabling more stable and sample-efficient policy extraction. IC-IQL (tarasov2025yes, ) integrates a Q-learning objective into AD. The policy network is updated by optimizing only its head, whereas the value function update propagates through both the head and the transformer backbone. In contrast, we train the whole multi-head transformer policy end-to-end, releasing the transformer’s scalability with a simplified pipeline. Also, we design a lightweight prompt structure to overcome the limitation of AD algorithms that require long training histories as context. Empirical results in Sec. 5 also demonstrate our superiority to these two baselines.
3 Preliminaries: In-Context Reinforcement Learning
We consider a multi-task offline RL setting, where tasks follow a distribution $M^{i}\!=\!\langle\mathcal{S},\!\mathcal{A},\!\mathcal{T}_{i},\mathcal{R}_{i},\gamma\rangle$ $\sim P(M)$. Tasks share the same state-action spaces $\mathcal{S},\mathcal{A}$, while differing in the reward function $\mathcal{R}$ or transition dynamics $\mathcal{T}$. An offline dataset $\mathcal{D}^{i}\!=\!\sum_{j}(s_{j}^{i},a_{j}^{i},r_{j}^{i},s_{j}^{i^{\prime}})$ is collected by arbitrary behavior policies for each task out of a total of $N$ training ones. The agent can only access the offline datasets $\{\mathcal{D}^{i}\}_{i=1}^{N}$ to train an in-context policy as $\pi_{\theta}(a^{i}|s^{i};\beta^{i})$, where $\beta^{i}$ is a prompt that encodes task-relevant information. $\beta$ can be the preceding learning history in AD and a sequence of transitions in DPT.
During testing, the trained policy is evaluated on unseen tasks sampled from $P(M)$ through direct interaction with the environment. With fixed policy parameters $\theta$, all adaptations happen via the prompt/context $\beta$ that is initially empty and gradually constructed from history interactions. As $\beta$ evolves, the model revises its predictions in a manner analogous to the policy updates in conventional RL algorithms. The objective is to learn an in-context policy that maximizes the expected episodic return over test tasks as $J(\pi)=\mathbb{E}_{M\sim P(M)}[J_{M}(\pi)]$.
@@ 4 Method
In this section, we present SICQL (Scalable In-Context Q-Learning), an innovative ICRL framework that leverages dynamic programming and world modeling for efficient reward maximization and task generalization. Figure 1 illustrates the method overview. The algorithm pseudocodes are presented in Appendix B, and detailed implementations are given as follows
@@
Figure 1: The overview of SICQL. (a) We pretrain a generalized world model to accurately capture task-relevant information from the multi-task offline dataset as in Eq. (2), and use the context encoder to transform a small number of raw transitions into a precise and lightweight prompt $\beta$ as in Eq. (3). (b) We design a prompt-based multi-head transformer model that simultaneously predicts the optimal policy $\pi_{\theta}(a|s;\beta)$, the state value function $V_{\theta}(s;\beta)$, and Q-function $Q_{\theta}(s,a;\beta)$ using separate heads, given the task prompt $\beta$ and corresponding query inputs ($s$ or $s,a$). We learn $V_{\theta}$ by expectile regression as in Eq. (5), and use it to compute Bellman backups for training $Q_{\theta}$ as in Eq. (4). The in-context value functions are distilled into policy extraction using advantage-weighted regression as in Eq. (6). (c) Online testing by interacting with the environment. The prompt is initially empty and gradually constructed from history interactions using the pretrained context encoder.
4.1 Prompt Construction via World Modeling
A key obstacle in RL generalization across tasks is how to accurately capture task-relevant information from offline datasets. In language communities, prompts consist of concise and precise text instructions rich in semantic information, as words inherently convey high-level concepts that represent entire semantic entities like objects (nouns) and actions (verbs) (kakogeorgiou2022hide, ). In contrast, ICRL approaches typically use a collection of raw state-action-reward transitions to infer the underlying task and identify what actions lead to good outcomes (lee2023supervised, ). The resulting prompt has many more tokens than a sentence and can be highly redundant. For instance, the states in a trajectory change smoothly with minor local variations (nagabandi2018neural, ), and in sparse-reward scenarios, only a very small fraction of transition data contains valuable information (burda2019exploration, ). Moreover, in offline RL settings, the task information can be entangled with behavior policies, leading to biased task inference at test time due to the change of behavior policies (yuan2022robust, ; wang2024meta, ). Overall, the current type of prompt may be insufficient to achieve precise in-context task inference for ICRL.
@@ 4.1 Prompt Construction via World Modeling
To address this challenge, we aim to design a lightweight prompt structure capable of encoding precise information about decision tasks. RL agents learn by actively interacting with the external environment, modeled as an MDP. The world model, representing environment dynamics (hafner2025mastering, ), fully characterizes the underlying task and remains invariant to behavior policies or the datasets collected. Inspired by this fundamental property of RL, we pretrain a generalized world model to acquire task-relevant information from the multi-task offline dataset, and use it to transform raw transitions $(s,a,r)$ into compact task representations $z$ to construct the prompt.

### v2

@@ Abstract
Recent advancements in language models have demonstrated remarkable in-context learning abilities, prompting the exploration of in-context reinforcement learning (ICRL) to extend the promise to decision domains. Due to involving more complex dynamics and temporal correlations, existing ICRL approaches may face challenges in learning from suboptimal trajectories and achieving precise in-context inference. In the paper, we propose Scalable In-Context Q-Learning (S-ICQL), an innovative framework that harnesses dynamic programming and world modeling to steer ICRL toward efficient reward maximization and task generalization, while retaining the scalability and stability of supervised pretraining. We design a prompt-based multi-head transformer architecture that simultaneously predicts optimal policies and in-context value functions using separate heads. We pretrain a generalized world model to capture task-relevant information, enabling the construction of a compact prompt that facilitates fast and precise in-context inference. During training, we perform iterative policy improvement by fitting a state value function to an upper-expectile of the Q-function, and distill the in-context value functions into policy extraction using advantage-weighted regression. Extensive experiments across a range of discrete and continuous environments show consistent performance gains over various types of baselines, especially when learning from suboptimal data. Our code is available at https://github.com/NJU-RL/SICQL.
@@ 1 Introduction
A longstanding goal of reinforcement learning (RL) is to learn from diverse experiences and generalize beyond its training environments, efficiently adapting to unseen situations, dynamics, or objectives (Ackley & Littman, 1989; Kirk et al., 2023). A promising avenue is in-context learning (Brown et al., 2020) that trains large-scale transformer models on massive datasets to achieve remarkable generalization capabilities, adapting to new tasks via prompt conditioning without any model updates (Wang et al., 2023; Li et al., 2024a). Accordingly, in-context RL (ICRL) seeks to extend this promise to decision domains and has seen rapid progress in recent years (Nikulin et al., 2025). Existing studies contain two typical branches: algorithm distillation (AD) (Laskin et al., 2023) and decision-pretrained transformer (DPT) (Lee et al., 2023), due to their simplicity and generality. They commonly employ cross-episode transitions as few-shot prompts and train transformer-based policies under supervised pretraining (Lin et al., 2024), followed by various improvements from model-based planning (Son et al., 2025), hierarchical decomposition (Huang et al., 2024), importance weighting (Dong et al., 2024), etc (Sinii et al., 2024; Tarasov et al., 2025; Dai et al., 2024).
@@
Though, significant challenges may emerge when extending the promise of in-context learning from (self-) supervised learning to RL, since RL involves more complex dynamics and temporal correlations (Silver et al., 2021). First, previous studies usually adopt a supervised pretraining paradigm, failing to go beyond imitating collected data (Yamagata et al., 2023). AD can require long-horizon context and inherit suboptimal behaviors due to the gradual update rule (Son et al., 2025). DPT relies on an oracle for optimal action labeling that can often be infeasible in practice (Tarasov et al., 2025). These limitations may hinder the efficient learning of optimal policies, especially when only suboptimal datasets are available. Second, substantial scope remains to advance the design of efficient prompts that can precisely encode RL task information. In language communities, prompts are concise, precise text instructions rich in semantic information, as text naturally conveys high-level concepts like objects (nouns) and actions (verbs) (Kakogeorgiou et al., 2022). However, ICRL approaches generally take raw transitions as prompts that have many more tokens than a sentence and can be highly redundant. Within transitions from offline datasets, the task information can be entangled with behavior policies, thus producing biased task inference at test time (Yuan & Lu, 2022). Hence, this type of prompt may be insufficient to precisely capture relevant information about decision tasks. The aforementioned limitations raise a key question: Can we design a scalable ICRL framework using lightweight prompts that precisely capture task-relevant information, while unleashing the core potential of fundamental reward maximization to learn from suboptimal data?
@@
To tackle these challenges, we draw upon two basic properties inherent in full RL. First, classical RL algorithms learn a value function to backpropagate expected returns using dynamic programming updates,  [^1] In this paper, we use dynamic programming to indicate the fundamental characteristic of any RL algorithm relying on the Bellman-backup operation. It updates state (or state-action) values based on value estimates of successor states (or state-action pairs), i.e., bootstrapping (Sutton et al., 1998). We use dynamic programming and Q-learning interchangeably to refer to this fundamental property of RL. showing appealing stitching property, i.e., the ability to combine parts of suboptimal trajectories for finding globally optimal behaviors (Sutton et al., 1998). Naturally, the stitching property offers a compelling avenue for unleashing the potential of ICRL architectures, attaining substantial improvement over suboptimal data. Second, as RL agents learn through active interactions with the outer environment, the decision task is fully characterized by the environment dynamics, i.e., the state transition and reward functions $p(s^{\prime},a|s,a)$. The world model (Hafner et al., 2025) can learn an internal representation of the environment dynamics, and is intrinsically invariant to behavior policies or collected datasets (Wang et al., 2024). Hence, leveraging the world model holds promise for designing a lightweight prompt structure capable of precisely encoding task-relevant information.
@@
Drawing inspiration from full RL, we propose Scalable In-Context Q-Learning (S-ICQL), an innovative framework that harnesses dynamic programming and world modeling to steer ICRL toward efficient reward maximization and task generalization. First, we design a prompt-based multi-head transformer architecture to maintain scalability and parameter efficiency. The model simultaneously predicts optimal policies and in-context value functions using separate heads, given a task prompt and corresponding query inputs. Second, we pretrain a generalized world model to capture task-relevant information from the multi-task offline dataset, and use it to transform a small number of raw transitions into a lightweight prompt for fast and precise in-context inference. Finally, we perform iterative policy improvement by fitting a state value function to an upper-expectile of the Q-function, and distill the in-context value functions into policy extraction using advantage-weighted regression. This formulation allows for learning a policy to maximize the Q-values subject to an offline dataset constraint, while retaining the scalability and stability of the supervised pretraining paradigm. In summary, our main contributions are threefold:
@@ 2 Related Work
The concept of agents adapting their behaviors within the context without model updates builds on earlier work in meta-RL, notably RL2 (Duan et al., 2016) that includes past interactions as inputs to an RNN-based policy. Recently, there has been a shift towards employing transformers to implement ICRL (Nikulin et al., 2025), driven by their proven ability to capture long-term dependencies and exhibit emergent in-context learning behaviors Brown et al. (2020). In online RL settings, AMAGO (Grigsby et al., 2024a; b) trains long-sequence transformers over entire rollouts with actor-critic learning to tackle in-context goal-conditioned problems, and (Lu et al., 2023) leverages the S4 (structured state space sequence) model’s ability to handle long-range sequences for ICRL tasks. In offline settings, two classical branches are AD (Laskin et al., 2023) and DPT (Lee et al., 2023). AD trains a causal transformer to autoregressively predict actions using preceding learning histories as context, while DPT predicts the optimal action based on a query state and a prompt of interaction transitions. Follow-up studies enhance in-context learning from different perspectives (Son et al., 2025; Huang et al., 2024; Sinii et al., 2024; Dai et al., 2024; Zisman et al., 2024). For example, IDT (Huang et al., 2024) designs a hierarchical decision structure, and DICP (Son et al., 2025) incorporates model-based planning with a learned dynamics model. These approaches generally adopt a supervised paradigm with raw transitions as prompts, while we harness dynamic programming for reward maximization and leverage world modeling to construct more efficient prompts.
@@
DIT (Dong et al., 2024) and IC-IQL (Tarasov et al., 2025) are the most relevant to our work. DIT also uses a weighted maximum likelihood estimation loss to train the transformer policy, where the weights are directly calculated from observed rewards in collected datasets. In contrast, we learn separate in-context value functions for computing advantage weights, enabling more stable and sample-efficient policy extraction. IC-IQL integrates a Q-learning objective into AD. The policy and value networks are updated by optimizing only their heads, without propagating gradients to the transformer backbone. In contrast, we train the full multi-head transformer policy end-to-end, releasing the transformer’s scalability with a simplified pipeline. We also design a precise, lightweight prompt structure to overcome the limitation of AD algorithms that require long training histories as context. Empirical results in Sec. 4 demonstrate our superiority to these two baselines.
@@ 3 Method
In this section, we present S-ICQL (Scalable In-Context Q-Learning), an innovative ICRL framework that leverages dynamic programming and world modeling for efficient reward maximization and task generalization. We adopt a prompt-based multi-head transformer architecture that simultaneously predicts optimal policy and in-context value functions using separate heads, ensuring scalability and parameter efficiency. Figure 1 illustrates the method overview. The algorithm pseudocodes are presented in Appendix B, and detailed implementations are given as follows.
@@
Figure 1: The overview of S-ICQL. (a) We pretrain a generalized world model to accurately capture task-relevant information from the multi-task offline dataset as in Eq. (3), and use the context encoder to transform a small number of raw transitions into a precise and lightweight prompt $\beta$ as in Eq. (1). (b) We design a prompt-based multi-head transformer model that simultaneously predicts the optimal policy $\pi_{\theta}(a|s;\beta)$, the state value function $V_{\theta}(s;\beta)$, and Q-function $Q_{\theta}(s,a;\beta)$ using separate heads, given the task prompt $\beta$ and corresponding query inputs ($s$ or $s,a$). We learn $V_{\theta}$ by expectile regression as in Eq. (5), and use it to compute Bellman backups for training $Q_{\theta}$ as in Eq. (4). The in-context value functions are distilled into policy extraction using advantage-weighted regression as in Eq. (6). (c) Online testing by interacting with the environment. The prompt is initially empty and gradually constructed from history interactions using the pretrained context encoder.
3.1 Problem Statement
We consider a multi-task offline RL setting, where tasks follow a distribution $M^{i}\!=\!\langle\mathcal{S},\!\mathcal{A},\!\mathcal{T}^{i},\mathcal{R}^{i},\gamma\rangle$ $\sim P(M)$. Tasks share the same state-action spaces $\mathcal{S},\mathcal{A}$ but differ in reward functions $\mathcal{R}$ or transition dynamics $\mathcal{T}$. An offline dataset $\mathcal{D}^{i}\!=\!\sum_{j}(s_{j}^{i},a_{j}^{i},r_{j}^{i},s_{j^{\prime}}^{i})$ is collected by arbitrary behavior policies for each task out of a total of $N$ training ones. The agent can only access the offline datasets $\{\mathcal{D}^{i}\}_{i=1}^{N}$ to train an in-context policy as $\pi_{\theta}(a^{i}|s^{i};\beta^{i})$, where $\beta^{i}$ is a prompt that encodes task-relevant information (e.g., past interaction history in AD or transition sequences in DPT). During testing, the trained policy is evaluated on unseen tasks sampled from $P(M)$ through direct interaction with the environment. With fixed policy parameters $\theta$, all adaptations occur through the prompt/context $\beta$ that is initially empty and gradually constructed from history interactions. As $\beta$ evolves, the model refines its predictions, analogous to policy updates in conventional RL. The objective is to learn an in-context policy that maximizes the expected episodic return over test tasks as $J(\pi)=\mathbb{E}_{M\sim P(M)}[J_{M}(\pi)]$.
3.2 Scalable Model Architecture
World Modeling.
@@ 3.2 Scalable Model Architecture
As shown in Figure 1-(a), we design a lightweight prompt structure capable of encoding precise information about decision tasks. The world model, representing environment dynamics $p(s^{\prime},r|s,a)$ (Hafner et al., 2025), fully characterizes the underlying task and remains invariant to behavior policies or the datasets collected. Inspired by this fundamental property of RL, we pretrain a generalized world model to acquire task-relevant information from the multi-task offline dataset. The world model contains a context encoder $E_{\phi}$ that abstracts recent $k$-step experiences $\eta_{t}^{i}\!=\!(s_{t-k},a_{t-k},r_{t-k},...,s_{t-1},a_{t-1},r_{t-1},s_{t},a_{t})^{i}$ into a task representation as $z^{i}_{t}\!=\!E_{\phi}(\eta_{t}^{i})$, and a dynamics decoder $D_{\varphi}$ that predict the instant reward and next state conditioned on the task representation as $\hat{r}_{t},\hat{s}_{t+1}]\!=\!D_{\varphi}(s_{t},a_{t};z_{t}^{i})$. Then, for training the in-context policy, we sample a short $h$-step trajectory randomly from dataset $\mathcal{D}^{i}$, and use the pretrained $E_{\phi}$ to transform raw transitions in this trajectory into compact task representations as
$\beta^{i}:=[z_{1}^{i},z_{2}^{i},...,z_{h}^{i}]=\left[E_{\phi}(\eta_{1}^{i}),E_{\phi}(\eta_{2}^{i}),...,E_{\phi}(\eta_{h}^{i})\right],~~~\forall~\text{task}~i.$ (1)
Then, we construct a lightweight prompt $\beta$ using the $h$-step task representation for fast and precise in-context inference, as opposed to AD algorithms that can require long training histories as context. Sec. 3.3 presents the world model and its pretraining process in detail.
Dynamic Programming.
Classical RL algorithms learn a value function to backpropagate expected returns using dynamic programming updates, showcasing high stitching capacity that allows for not only imitating collected data but also achieving substantial improvement beyond it (Sutton et al., 1998). Inspired by this, we harness the stitching property to offer a promising avenue for unleashing ICRL’s potential toward explicit reward maximization. Following common practice in advanced offline RL algorithms (Kostrikov et al., 2022; Snell et al., 2023), we learn both a state value function $V$ and an action value function $Q$. As shown in Figure 1-(b), we design a prompt-based multi-head transformer architecture to simultaneously predict the optimal and value functions. Let $\theta$ denote parameters of the integrated transformer model. The model outputs the policy $\pi_{\theta}(a|s;\beta)$, the state value $V_{\theta}(s;\beta)$, and the action value $Q_{\theta}(s,a;\beta)$ using three separate heads, given the task prompt $\beta$ and corresponding query inputs ($s$ or $s,a$). Sec. 3.4 presents the detailed learning process.
This flexible design maintains two appealing properties. One is parameter efficiency. Our model only introduces two additional lightweight heads compared to traditional ICRL methods such as DPT, and the resulting increase in parameters is negligible relative to the foundational transformer backbone. Another is model scalability. Our central component remains an end-to-end causal transformer architecture that holds the promise for training RL models at scale. We learn the in-context policy by advantage-weighted regression, a supervised learning style with a simple and convergent maximum likelihood loss (refer to Eq. (6)). It ensures that our method preserves the scalability and stability inherent to the supervised pretraining paradigm. Moreover, we train the whole multi-head transformer policy end-to-end (refer to Sec. 3.5), releasing the transformer’s scalability with a simplified pipeline.
3.3 World Model Pretraining for Prompt Construction